# Chain of Code: Reasoning with a Language Model-Augmented Code Emulator

### v2

@@ Abstract
https://chain-of-code.github.io
Code provides a general syntactic structure to build complex programs and perform precise computations when paired with a code interpreter – we hypothesize that language models (LMs) can leverage code-writing to improve Chain of Thought reasoning not only for logic and arithmetic tasks [5, 26, 1], but also for semantic ones (and in particular, those that are a mix of both). For example, consider prompting an LM to write code that counts the number of times it detects sarcasm in an essay: the LM may struggle to write an implementation for “detect_sarcasm(string)” that can be executed by the interpreter (handling the edge cases would be insurmountable). However, LMs may still produce a valid solution if they not only write code, but also selectively “emulate” the interpreter by generating the expected output of “detect_sarcasm(string)” and other lines of code that cannot be executed. In this work, we propose Chain of Code (CoC), a simple yet surprisingly effective extension that improves LM code-driven reasoning. The key idea is to encourage LMs to format semantic sub-tasks in a program as flexible pseudocode that the interpreter can explicitly catch undefined behaviors and hand off to simulate with an LM (as an “LMulator"). Experiments demonstrate that Chain of Code outperforms Chain of Thought and other baselines across a variety of benchmarks; on BIG-Bench Hard, Chain of Code achieves 84%, a gain of 12% over Chain of Thought. CoC scales well with large and small models alike, and broadens the scope of reasoning questions that LMs can correctly answer by “thinking in code".
@@
Chain of Thought
Chain of Code
@@
Figure 1: Chain of Code generates code and reasons through an LM-augmented code emulator. Lines evaluated with Python are in red and with an LM are in purple. The full query is in Fig. LABEL:fig:intro_query. (0(a)-0(c)) show results on BIG-Bench Hard compared to human performance [34].
@@ 1 Introduction
Language models (LMs) at certain scale exhibit the profound ability to solve complex reasoning questions [3, 41] – from writing math programs [9] to solving science problems [17]. Notably, these capabilities have shown to improve with Chain of Thought (CoT) prompting [42], whereby complex problems are decomposed into a sequence of intermediate reasoning steps. CoT excels at semantic reasoning tasks, but tends to struggle with questions that involve numeric or symbolic reasoning [36, 23]. Subsequent work addresses this by prompting LMs (e.g., trained on Github [4]) to write and execute code [5, 26, 1]. Code in particular is advantageous because it provides both (i) a general syntactic structure to build and encode complex programs [19] (e.g., logic structures, functional vocabularies – in ways that are Turing complete), and (ii) an interface by which existing APIs paired together with an interpreter can be used to perform precise algorithmic computations (e.g., from multiplication of large numbers to sorting an array of size 10,000) that a language model trained only to mimic the statistically most likely next token would otherwise struggle to produce.
@@
While writing and executing code may improve LM reasoning performance across a wide range of arithmetic tasks, this particular approach contends with the fact that many semantic tasks are rather difficult (and at times, nearly impossible) to express in code. For example, it remains unclear how to write a function that returns a boolean when it detects sarcasm in a string [36] (handling the edge cases would be insurmountable). Perhaps fundamentally, using LMs to write programs in lieu of multi-step textual reasoning inherently assumes that the intermediate reasoning traces (expressed in lines of code) all need to be executable by an interpreter. Is it possible to lift these restrictions to get the best of both reasoning in code and reasoning in language?
@@
In this work, we propose Chain of Code (CoC), a simple yet surprisingly effective extension to improve LM code-driven reasoning – where the LM not only writes a program, but also selectively “simulates” the interpreter by generating the expected output of certain lines of code (that the interpreter could not execute). The key idea is to encourage LMs to format semantic sub-tasks in a program as flexible pseudocode that at runtime can be explicitly caught and handed off to emulate with an LM – we term this an LMulator (a portmanteau of LM and emulator). For example, given the task “in the above paragraph, count how many times the person was sarcastic,” we can in-context prompt the LM to write a program that may call helper functions such as is_sarcastic(sentence), to which the LM makes a linguistic prediction and returns the result as a boolean output, that then gets processed with the rest of the program. Specifically, we formulate LM reasoning as the following process (illustrated in Figure 1): the LM writes code, the interpreter steps through to execute each line of code (in red), or if it fails, simulates the result with the LM (in purple) and updates the program state (in green). CoC inherits the benefits of both (i) writing executable code (where precise algorithmic compututations are left to an interpreter), and (ii) writing pseudocode for semantic problems, and generating their outputs (which can be thought of as a simple formatting change, to which LMs are robust [22]) – enabling the LM to “think in code.”
@@
Extensive experiments demonstrate that CoC is applicable to a wide variety of challenging numerical and semantic reasoning questions, and outperforms a number of popular baselines. In particular, we find that it achieves high performance on BIG-Bench Hard tasks [36], outperforming average human raters overall and even the best human raters on an algorithmic subset of tasks, and to the best of our knowledge setting a new state of the art. We further show that both code interpreter execution and language model execution simulation are necessary for this performance, and that the approach scales well with large and small models alike – contrary to prompting techniques like Chain of Thought that only emerge at scale. Finally, we demonstrate how Chain of Code can serve as a general purpose reasoner via cross-task prompting benchmark, which in contrast to prior work, uses prompts from different families of problems as context – providing only the structure of the response (as opposed to the solution itself). This work underscores how one may leverage the structure and computational power of code and the reasoning abilities of language models to enable a “best of both worlds” reasoner.
@@ 2 Chain of Code: Reasoning with an LMulator
In this section, we describe Chain of Code (CoC) prompting, an approach that leverages the ability of language models to code, to reason, and to leverage an LM-augmented code emulator (an LMulator) to simulate running code. We start with background in Section 2.1, then overview the method in Section 2.2, its implementation in Section 2.3, and finally its capabilities in Section 2.4.
@@ 2.1 Preliminaries
Briefly, we overview some background on LM reasoning. Many of these reasoning techniques have been enabled by in-context learning [3], which provides the model with a few demonstrative examples at inference time, rather than updating any weights with gradients. These examples serve to provide context and format for the setting, enabling the model to emulate these examples while adapting to a new query. This property has been instrumental in easily applying LMs to new tasks as it can be rapidly adapted and requires minimal data.
Through in-context learning, approaches have been developed to leverage human thought processes and use tools to improve performance of language models. We outline three such approaches that provide the foundations for Chain of Code. Chain of Thought (CoT) [42], ScratchPad [26], and Program of Thoughts [5] demonstrated the efficacy of breaking problems down into substeps. For CoT these substeps are in natural language, mirroring one’s thought process when stepping through a complicated problem. ScratchPad, on the other hand, maintains a program state of intermediate steps when simulating the output of code – resulting in an LM acting as a code interpreter. Program of Thoughts [5] focused on generating the code itself, which is then executed by a code interpreter to solve reasoning problems. Each of these is visualized in Figure 1(c).
(a) Chain of Thought [42]
(b) Program of Thoughts [5]
(c) ScratchPad [26]
(d) Chain of Code Generation
(e) Chain of Code Execution
Figure 2: Previous reasoning methods: To solve advanced problems, (1(a)) Chain of Thought prompting breaks the problem down into intermediate steps, (1(b)) Program of Thoughts prompting writes and executes code, and (1(c)) ScratchPad prompting simulates running already written code by tracking intermediate steps through a program state. Our reasoning method: Chain of Code first (LABEL:fig:method_generation) generates code or psuedocode to solve the question and then (LABEL:fig:method_execution) executes the code with a code interpreter if possible, and with an LMulator (language model emulating code) otherwise. Blue highlight indicates LM generation, red highlight indicates LM generated code being executed, and purple highlight indicates LMulator simulating the code via a program state in green.
@@ 2.2 Chain of Code
Inspired by how a human may reason through a particularly complex problem with a mix of natural language, pseudocode, and running code or how a researcher may develop a new general algorithm through a code-based formalism then apply it to a problem, Chain of Code proceeds in two steps: (1) Generation, which, given the question to solve, an LM generates code to reason through the problem, and (2) Execution, which executes the code via a code interpreter when possible and via an LM when not. See Section 2.3 for more details on the specific implementation.
@@ 2.3 Chain of Code Implementation
As a brief example, the code answer = 0; answer += is_sarcastic(‘you don’t say’); answer += 1; would be executed as follows: (1) Python would execute the first line answer = 0; and update the program state to {answer = 0}, (2) Python would attempt to execute the second line and fail, and thus the LMulator would simulate the code answer += is_sarcastic(‘you don’t say’); by generating the program state {answer = 1}, which would be updated in the program, (3) Python would execute the last line answer += 1; and update the program state to {answer = 2}, (4) the answer would be retrieved as 2.
@@ 2.4 Chain of Code Abilities
3 Language Reasoning Experimental Evaluation
@@
    How well does CoC perform overall across a variety of tasks?
@@
    How does each aspect of CoC affects overall performance?
@@
    How does CoC compare with instruction tuned chat models with and without tools?
@@ 3.1 Baselines and Ablations
We consider our main method to be CoC (Interweave), also referred to as CoC (Ours), though we also propose two variants with simpler implementation and modestly lower performance: CoC (try Python except LM) and CoC (try Python except LM state). These two variants attempt to run the entire generated code with Python (rather than line by line) and if it fails, simulate the code execution with the LMulator, outputting a final answer or an intermediate state trace, respectively. We also perform the following ablations, some of which are comparable to previous work as noted. In CoC (Python) Python is used to run the entire generated code and if the code is not executable, it is marked as failure – this can be thought of as a comparison to Program of Thoughts [5] or Program-aided language models [10]. We note that in many cases this baseline is particularly challenged, as writing executable code for some of the reasoning problems becomes nearly impossible (e.g., writing code to judge if a phrase is sarcastic), but one may focus on the results for Algorithmic only tasks for a more fair comparison. In CoC (LM) the code is interpreted by an LMulator outputting the final answer, and in CoC (LM state) the code is interpreted by an LMulator outputting a state trace of intermediate steps – this can be thought of as ScratchPad prompting for reasoning [26]. Note, the last two ablations do not leverage the Python interpreter.
We also compare against the following baselines. In Direct question answering the LM simply responds to the question with a final answer. In Chain of Thought prompting (CoT) the LM uses intermediate steps to solve the task; we use CoT as our standard prompt technique for the field of substep prompting [16, 48] as prompts are readily available.
@@ 3.2 Tasks
We consider a subset of challenging tasks from BIG-Bench [34] called BIG-Bench Hard (BBH) [36] to ensure we are solving the most challenging tasks. These tasks were specifically selected for their difficulty for language models and the datasets provides human-rater baselines and a set of Chain of Thought prompts. The 23 tasks require semantic reasoning(e.g., “Movie Recommendation”), numerical reasoning (e.g., “Multi-Step Arithmetic”), and a combination of both (e.g., “Object Counting”). As such they enable us to study the efficacy of CoC across varied problems, not just those that coding is a natural fit for. Several prompts are shown in Appendix Figure A1. We also show results for the grade-school math (GSM8K) benchmark [7] in Appendix Section A.2, though find that these problems are primarily solved algorithmically alone through code.
@@
The models used herein include the OpenAI family of models: text-ada-001, text-baggage-001, text-curie-001, and text-davinci-003 (in plots we denote these as a-1, b-1, c-1, and d-3). We also consider PaLM-2’s code finetuned variant [6, 12]. For instruction tuned models, we compare to recent variants of GPT (gpt-3.5-turbo and gpt-4) with the chat completion mode run in October 2023. The results below are using the text-davinci-003 model unless otherwise stated.
@@ 3.3 Results
Question 1: Overall Performance. The overall performance of CoC is shown in Figure 1 and Table 1 (with full results in Table A1). We see that CoC outperforms other approaches, both in the number of tasks it exceeds the human baseline and in the overall amount that it exceeds the baseline. Indeed, CoC’s 84% is SoTA to the best of our knowledge [11]. In several tasks CoC vastly outperforms the human baseline and other methods, achieving nearly 100% – generally for these tasks the result is complicated in language but trivial in code (e.g., a task from multi-step arithmetic Q: ((3+5×8×4)(98×7))=((-3+5$\times$ 8$\times$-4)-(9-8$\times$-7))=( ( - 3 + 5 × 8 × - 4 ) - ( 9 - 8 × - 7 ) ) =). We also observe that CoT outperforms the human baseline on a number of tasks, while the Direct answer fares poorly.
Table 1: Overall performance (%) with both few-shot prompting with a single task and cross-task. The delta compared to direct prompting is shown in parenthesis.
text-davinci-003 PaLM 2-S* (code variant [12])
Human Direct CoT CoC Direct CoT CoC
@@
Question 2: Problem Type. Figure 3 breaks the results down by problem type; the task labels are shown in Table A1. First, we isolate problems that are primarily algorithmic or primarily natural language (these categories were identified in [36]). We see that on algorithmic tasks, CoC performs particularly well, while on natural language tasks CoC performs on par with CoT. This is particularly encouraging, because one may expect these language oriented tasks to be a worse fit for code. The key is that our method offers the flexibility of using a LMulator to simulate the output of code execution, retaining the semantic reasoning capabilities of LMs for natural language problems.
@@
Figure 3: Average performance across different baselines grouped by task type, indicating the problem type and how the CoC is generated and executed.
Question 3: Ablations. Figures 4 and  5, and Table 2 show the ablations performed to motivate each aspect of Chain of Code prompting. As one may expect, the approaches that execute Python (CoC (Interweave, Python, try Python except LM, try Python except LM state)) achieve 100% performance on several tasks – if the code is correct, then the model will be correct every time. However, the approach that relies on Python alone (CoC (Python)) performs poorly when applied to non-algorithmic tasks, failing almost all. The CoC (Python) ablation is similar to recent works [10, 5], which show that if applied to numerical problems then code reasoning performs well. CoC without the Python interpreter (CoC (LM, LM state)) too fares poorly, though we see that the step-by-step approach proposed in ScratchPad prompting [26] improves in each task.
We also show that ablations CoC (try Python except LM, try Python except LM state), in which CoC first tries to run the entire code with Python and if it fails simulates the code with an LM, perform quite well. Again we see that maintaining a program state provides an improvement in performance. With only minor degradations in performance observed, they are reasonable alternatives to the fully interweaved CoC for their simplicity. Though we note, these ablations’ performance would be much worse in cases where interweaving code and semantics is truly necessary – for example, if we imagine a case where code is necessary to parse image inputs or to access an external database, but language is necessary to parse the results (see the robotics applications in Section 4).
Figure 4: CoC ablations on average performance grouped by task type.
Figure 5: Results across all BIG-Bench Hard tasks compared to human baseline [34]. The tasks (x-axis) in each plot are sorted individually by performance. See Table A1 and Figure 4 for a breakdown by task type.

### v3

@@ Abstract
Code provides a general syntactic structure to build complex programs and perform precise computations when paired with a code interpreter - we hypothesize that language models (LMs) can leverage code-writing to improve Chain of Thought reasoning not only for logic and arithmetic tasks (Chen et al., 2022; Nye et al., 2021; Austin et al., 2021), but also for semantic ones (and in particular, those that are a mix of both). For example, consider prompting an LM to write code that counts the number of times it detects sarcasm in an essay: the LM may struggle to write an implementation for “detect_sarcasm(string)” that can be executed by the interpreter (handling the edge cases would be insurmountable). However, LMs may still produce a valid solution if they not only write code, but also selectively “emulate” the interpreter by generating the expected output of “detect_sarcasm(string)”. In this work, we propose Chain of Code (CoC), a simple yet surprisingly effective extension that improves LM code-driven reasoning. The key idea is to encourage LMs to format semantic sub-tasks in a program as flexible pseudocode that the interpreter can explicitly catch undefined behaviors and hand off to simulate with an LM (as an “LMulator"). Experiments demonstrate that Chain of Code outperforms Chain of Thought and other baselines across a variety of benchmarks; on BIG-Bench Hard, Chain of Code achieves 84%, a gain of 12% over Chain of Thought. In a nutshell, CoC broadens the scope of reasoning questions that LMs can answer by “thinking in code".
Machine Learning, ICML
{CJK*}
UTF8gbsn
@@
Q: How many countries have I been to? I’ve been to Mumbai, London, Washington, Grand Canyon, ...
A: 32 (20%, ), 29 (10%, ), 52 (10%, ), ...
@@
Q: Let’s think step by step. How many countries have I been to? I’ve been to Mumbai, London, ...
We’ll group by countries and count: 1. India: Mumbai, Delhi, Agra 2. UK: London, Dover, Edinburgh, Skye 3. USA: Washington, Grand Canyon, ... A: 61 (20%, ), 60 (20%, ), 52 (10%, ), ...
Chain of Code (Ours)
Q: How many countries have I been to? I’ve been to Mumbai, London, Washington, Grand Canyon, Baltimore, ...
1    places, countries = ["Mumbai", ...], set() delta state: {places = [‘Mumbai’, ...], countries = set()}
2    for place in places: delta state: {place = ‘Mumbai’}
3      country = get_country(place) delta state: {country = ‘India’)}
4      countries.add(country) delta state: {countries = {‘India’}}
5    answer = len(countries)    delta state: {answer = 52}
A: 52 (100%, )
Figure 1: Chain of Code generates code and reasons through an LM-augmented code emulator. Lines evaluated with Python are in red and with an LM are in purple. The full query is in Fig. LABEL:fig:intro_query.
@@
Figure 2: Overall results on BIG-Bench Hard compared to human performance (Srivastava et al., 2022).
@@ 1 Introduction
Language models (LMs) at certain scale exhibit the profound ability to solve complex reasoning questions (Brown et al., 2020; Wei et al., 2022a) – from writing math programs (Drori et al., 2022) to solving science problems (Lewkowycz et al., 2022). Notably, these capabilities have shown to improve with Chain of Thought (CoT) prompting (Wei et al., 2022b), whereby complex problems are decomposed into a sequence of intermediate reasoning steps. CoT excels at semantic reasoning tasks, but tends to struggle with questions that involve numeric or symbolic reasoning (Suzgun et al., 2022; Mirchandani et al., 2023). Subsequent work addresses this by prompting LMs (e.g., trained on Github (Chen et al., 2021)) to write and execute code (Chen et al., 2022; Nye et al., 2021; Austin et al., 2021). Code in particular is advantageous because it provides both (i) a general syntactic structure to build and encode complex programs (Liang et al., 2023) (e.g., logic structures, functional vocabularies – in ways that are Turing complete), and (ii) an interface by which existing APIs paired together with an interpreter can be used to perform precise algorithmic computations (e.g., from multiplication of large numbers to sorting an array of size 10,000) that a language model trained only to mimic the statistically most likely next token would otherwise struggle to produce.
@@
While writing and executing code may improve LM reasoning performance across a wide range of arithmetic tasks, this particular approach contends with the fact that many semantic tasks are rather difficult (and at times, nearly impossible) to express in code. For example, it remains unclear how to write a function that returns a boolean when it detects sarcasm in a string (Suzgun et al., 2022) (handling the edge cases would be insurmountable). Perhaps fundamentally, using LMs to write programs in lieu of multi-step textual reasoning inherently assumes that the intermediate reasoning traces (expressed in lines of code) all need to be executable by an interpreter. Is it possible to lift these restrictions to get the best of both reasoning in code and reasoning in language?
@@
In this work, we propose Chain of Code (CoC), a simple yet surprisingly effective extension to improve LM code-driven reasoning – where the LM not only writes a program, but also selectively “simulates” the interpreter by generating the expected output of certain lines of code (that the interpreter could not execute). The key idea is to encourage LMs to format semantic sub-tasks in a program as flexible pseudocode that at runtime can be explicitly caught and handed off to emulate with an LM – we term this an LMulator (a portmanteau of LM and emulator). For example, given the task “in the above paragraph, count how many times the person was sarcastic,” we can in-context prompt the LM to write a program that may call helper functions such as is_sarcastic(sentence), to which the LM makes a linguistic prediction and returns the result as a boolean output, that then gets processed with the rest of the program. Specifically, we formulate LM reasoning as the following process (illustrated in Figure 1): the LM writes code, the interpreter steps through to execute each line of code (in red), or if it fails, simulates the result with the LM (in purple) and updates the program state (in green). CoC inherits the benefits of both (i) writing executable code (where precise algorithmic compututations are left to an interpreter), and (ii) writing pseudocode for semantic problems, and generating their outputs (which can be thought of as a simple formatting change, to which LMs are robust (Min et al., 2022)) – enabling the LM to “think in code”.
@@
Extensive experiments demonstrate that CoC is applicable to a wide variety of challenging numerical and semantic reasoning questions, and outperforms a number of popular baselines. In particular, we find that it achieves high performance on BIG-Bench Hard tasks (Suzgun et al., 2022), outperforming average human raters overall and outperforming even the best human raters on an algorithmic subset of tasks, and to the best of our knowledge setting a new state of the art. We further show that both code interpreter execution and language model execution simulation are necessary for this performance, and that the approach scales well with large and small models alike – contrary to prompting techniques like Chain of Thought that only emerge at scale. We then demonstrate how Chain of Code can serve as a general purpose reasoner via cross-task prompting benchmark, which in contrast to prior work, uses prompts from different families of problems as context – providing only the structure of the response (as opposed to the solution itself). Finally, we show CoC is complementary to more advanced instruction tuned chat models, robust against prompt variation, and applicable beyond language reasoning domain like robotics. This work underscores how one may leverage the structure and computational power of code and the reasoning abilities of language models to enable a “best of both worlds” reasoner.
@@ 2 Chain of Code: Reasoning with an LMulator
In this section, we describe Chain of Code (CoC), an approach that leverages the ability of language models to code, to reason, and to leverage an LM-augmented code emulator (an LMulator) to simulate running code. We start with background in Section 2.1, then overview the method in Section 2.2, its implementation in Section 2.3, and finally its capabilities in Section 2.4.
@@ 2.1 Preliminaries
Briefly, we overview some background on LM reasoning. Many of these reasoning techniques have been enabled by in-context learning (Brown et al., 2020), which provides the model with a few demonstrative examples at inference time, rather than updating any weights with gradients. These examples serve to provide context and format for the setting, enabling the model to emulate these examples while adapting to a new query. This property has been instrumental in easily applying LMs to new tasks as it can be rapidly adapted and requires minimal data.
Through in-context learning, approaches have been developed to leverage human thought processes and use tools to improve performance of language models. We outline three such approaches that provide the foundations for Chain of Code. Chain of Thought (CoT) (Wei et al., 2022b), ScratchPad (Nye et al., 2021), and Program of Thoughts (Chen et al., 2022) demonstrated the efficacy of breaking problems down into substeps. For CoT these substeps are in natural language, mirroring one’s thought process when stepping through a complicated problem. ScratchPad, on the other hand, maintains a program state of intermediate steps when simulating the output of code – resulting in an LM acting as a code interpreter. Program of Thoughts (Chen et al., 2022) focused on generating the code itself, which is then executed by a code interpreter to solve reasoning problems. Each of these is visualized in Figure 3(c).
(a) Chain of Thought
(b) Program of Thoughts
(c) ScratchPad
(d) Chain of Code Generation (Ours)
(e) Chain of Code Execution (Ours)
Figure 3: Previous reasoning methods: To solve advanced problems, (LABEL:fig:prelim-cot) Chain of Thought prompting breaks the problem down into intermediate steps, (LABEL:fig:prelim-pot) Program of Thoughts prompting writes and executes code, and (LABEL:fig:prelim-scratchpad) ScratchPad prompting simulates running already written code by tracking intermediate steps through a program state. Our reasoning method: Chain of Code first (LABEL:fig:method_generation) generates code or psuedocode to solve the question and then (LABEL:fig:method_execution) executes the code with a code interpreter if possible, and with an LMulator (language model emulating code) otherwise. Blue highlight indicates LM generation, red highlight indicates LM generated code being executed, and purple highlight indicates LMulator simulating the code via a program state in green.
@@ 2.2 Chain of Code
Inspired by how a human may reason through a particularly complex problem with a mix of natural language, pseudocode, and runnable code or how a researcher may develop a new general algorithm through a code-based formalism then apply it to a problem, Chain of Code proceeds in two steps: (1) Generation, which, given the question to solve, an LM generates code to reason through the problem, and (2) Execution, which executes the code via a code interpreter when possible and via an LM when not. See Section 2.3 for more details on the specific implementation.
@@ 2.3 Chain of Code Implementation
To illustrate with a brief example, the code answer = 0; answer += is_sarcastic(‘you don’t say’); answer += 1; would be executed as follows: (1) Python would execute the first line answer = 0; and update the program state to {answer = 0}, (2) Python would attempt to execute the second line and fail, and thus the LMulator would simulate the code answer += is_sarcastic(‘you don’t say’); by generating the program state {answer = 1}, which would be updated in the program, (3) Python would execute the last line answer += 1; and update the program state to {answer = 2}, (4) the answer would be retrieved as 2.
@@ 2.4 Chain of Code Abilities
3 Experimental Evaluation
@@
    How well does CoC perform across a variety of tasks?
@@
    How does each aspect of CoC affect performance?
@@
    How can CoC be used with instruction tuned chat models?
  7. 7.
    How robust CoC is against prompt variation?
  8. 8.
    Can CoC be applied beyond language reasoning tasks?
@@ 3.1 Baselines and Ablations
We consider our main method to be CoC (Interweave), also referred to as CoC (Ours), though we also propose two variants with simpler implementation and modestly lower performance: CoC (try Python except LM) and CoC (try Python except LM state). These two variants attempt to run the entire generated code with Python (rather than line by line) and if it fails, simulate the code execution with the LMulator, outputting a final answer or an intermediate state trace, respectively. We also perform the following ablations, some of which are comparable to previous work as noted. In CoC (Python) Python is used to run the entire generated code and if the code is not executable, it is marked as failure – this can be thought of as a comparison to Program of Thoughts (Chen et al., 2022) or Program-aided language models (Gao et al., 2023). We note that in many cases this baseline is particularly challenged, as writing executable code for some of the reasoning problems becomes nearly impossible (e.g., writing code to judge if a phrase is sarcastic), but one may focus on the results for Algorithmic only tasks for a more fair comparison. In CoC (LM) the code is interpreted by an LMulator outputting the final answer, and in CoC (LM state) the code is interpreted by an LMulator outputting a state trace of intermediate steps – this can be thought of as ScratchPad prompting for reasoning (Nye et al., 2021). Note, the last two ablations do not leverage the Python interpreter.
We also compare against the following baselines. In Direct question answering the LM simply responds to the question with a final answer. In Chain of Thought prompting (CoT) the LM uses intermediate steps to solve the task; we use CoT as our standard prompt technique for the field of substep prompting (Kojima et al., 2022; Zhou et al., 2022a) as prompts are readily available.
@@ 3.2 Tasks
We consider a subset of challenging tasks from BIG-Bench (Srivastava et al., 2022) called BIG-Bench Hard (BBH) (Suzgun et al., 2022) to ensure we are solving the most challenging tasks. These tasks were specifically selected for their difficulty for language models and the datasets provides human-rater baselines and a set of Chain of Thought prompts. The 23 tasks require semantic reasoning (e.g., “Movie Recommendation”), numerical reasoning (e.g., “Multi-Step Arithmetic”), and a combination of both (e.g., “Object Counting”). As such they enable us to study the efficacy of CoC across varied problems, not just those that coding is a natural fit for. Several prompts are shown in Figure A1. We also show results for the grade-school math (GSM8K) benchmark (Cobbe et al., 2021) in Section A.2, although we find that these problems are primarily solved algorithmically alone through code.
@@
The models used herein include the OpenAI family of models: text-ada-001, text-baggage-001, text-curie-001, and text-davinci-003 (in plots we denote these as a-1, b-1, c-1, and d-3). We also consider PaLM-2’s code finetuned variant (Chowdhery et al., 2022; Google et al., 2023). For instruction tuned models, we compare to recent variants of GPT (gpt-3.5-turbo and gpt-4) with the chat completion mode run in October 2023 and January 2024. The results below are using the text-davinci-003 model unless otherwise stated.
@@ 3.3 Results
Question 1: Overall Performance. The overall performance of CoC is shown in Figure 2 and Table 1 (with full results in Table A1). We see that CoC outperforms other approaches, both in the number of tasks it exceeds the human baseline and in the overall amount that it exceeds the baseline. Indeed, CoC’s 84% is SoTA to the best of our knowledge (Gemini Team, 2023). In fact, when combined with gpt-4, CoC achieves 91% (see Table A4). In several tasks CoC vastly outperforms the human baseline and other methods, achieving nearly 100% – generally for these tasks the result is complicated in language but trivial in code (e.g., a task from multi-step arithmetic Q: ((3+5×8×4)(98×7))=((-3+5$\times$ 8$\times$-4)-(9-8$\times$-7))=( ( - 3 + 5 × 8 × - 4 ) - ( 9 - 8 × - 7 ) ) =). We also observe that CoT outperforms the human baseline on a number of tasks, while the Direct answer fares poorly.
Table 1: Overall performance (%) on BIG-Bench Hard with both few-shot prompting with a single task and cross-task. The delta compared to direct prompting is shown in parenthesis.
text-davinci-003 PaLM 2-S* (code variant (Google et al., 2023))
Prompt Human Direct CoT CoC (Ours) Direct CoT CoC (Ours)
@@
Question 2: Problem Type. Figure 4 breaks the results down by problem type; the task labels are shown in Table A1. First, we isolate problems that are primarily algorithmic or primarily natural language (these categories were identified in (Suzgun et al., 2022)). We see that on algorithmic tasks, CoC performs particularly well, while on natural language tasks CoC performs on par with CoT. This is particularly encouraging, because one may expect these language oriented tasks to be a worse fit for code. The key is that our method offers the flexibility of using a LMulator to simulate the output of code execution, retaining the semantic reasoning capabilities of LMs for natural language problems.
@@
Figure 4: Average performance across different baselines grouped by task type, indicating the problem type and how CoC is generated & executed.
Question 3: Ablations. Figures 5 and 6, and Table 2 show the ablations performed to motivate each aspect of Chain of Code prompting. As one may expect, the approaches that execute Python (CoC (Interweave, Python, try Python except LM, try Python except LM state)) achieve 100% performance on several tasks – if the code is correct, then the model will be correct every time. However, the approach that relies on Python alone (CoC (Python)) performs poorly when applied to non-algorithmic tasks, failing almost all. The CoC (Python) ablation is similar to recent works (Gao et al., 2023; Chen et al., 2022), which show that if applied to numerical problems then code reasoning performs well. CoC without the Python interpreter (CoC (LM, LM state)) too fares poorly, though we see that the step-by-step approach proposed in ScratchPad prompting (Nye et al., 2021) improves in each task.
We also show that ablations CoC (try Python except LM, try Python except LM state), in which CoC first tries to run the entire code with Python and if it fails simulates the code with an LM, perform quite well. Again we see that maintaining a program state provides an improvement in performance. With only minor degradations in performance observed, they are reasonable alternatives to the fully interweaved CoC for their simplicity. Though we note, these ablations’ performance would be much worse in cases where interweaving code and semantics is truly necessary – for example, if we imagine a case where code is necessary to parse image inputs or to access an external database, but language is necessary to parse the results (see the robotics applications in Section A.6).
Figure 5: Chain of Code ablations on average performance grouped by task type.
Figure 6: Results across all BIG-Bench Hard tasks compared to human baseline (Srivastava et al., 2022). The tasks (x-axis) in each plot are sorted individually by performance. See Table A1 and Figure 5 for a breakdown by task type.